# Yunrui Li,a Hao Xu,a and Pengyu Honga

### v2

@@ 3.4 Performance by molecular weights
Fig.  5: Model performance comparison between small, medium, and large molecules by molecular weight. The model performs equally well between small and medium molecules, with a marginally reduced precision for larger molecules.
@@
3.5 Performance on saccharides
Fig.  7: Exemplary demonstration of our model’s performance on saccharides. The observed experimental signals and the predicted signals are colored in blue and orange, respectively. The prediction error (MAE) is shown in the bottom right corner within each plot.
@@ 3.5 Performance on saccharides
Saccharides, or carbohydrates, play critical roles in various biological processes involved in energy source and storage, cell signaling, cell adhesion, cell recognition, structural integrity of cells and tissues, as well as cognitive functions and metabolic regulation.35, 36, 37, 38 Despite their importance, elucidating the structures of saccharides is challenging due to their inherent structural complexity and diversity. This complexity arises from the diverse arrangements of monosaccharide units, varied anomeric configurations, and variable glycosidic linkages. Additionally, carbohydrates often lack the crystallinity required for high-resolution X-ray diffraction, unlike the well-defined crystalline structures of small molecules or proteins. Consequently, NMR spectroscopy, particularly through techniques such as HSQC, has emerged as an indispensable tool in unraveling the detailed structures of carbohydrates.39, 40 Forecasting HSQC cross peaks and aligning them with experimental data can assist in comprehending saccharide connectivity and stereochemistry, thus aiding in structural determination.
@@
Our model demonstrates excellent performance in predicting HSQC cross peaks for saccharides molecules, yielding an MAE of 1.95 for ${}^{13}\text{C}$ shifts and 0.17 for ${}^{1}\text{H}$ shifts (see Figure 6). This level of accuracy, consistent with the overall model performance in Table 1, demonstrates the model’s robustness in handling complex saccharide structures. Figure 7 shows the performance of our model on a few exemplar saccharides. These saccharides feature multiple ring structures and numerous stereogenic centers, contributing to the intricate nature of their HSQC spectra. Despite these inherent complexities, our model exhibits exceptional accuracy in predicting the HSQC cross peaks for these molecules. This robust performance underscores our model’s capacity to navigate the complexities associated with saccharides, thereby emphasizing its versatility and effectiveness across various applications in the field.
3.6 Effects of pre-training and fine-tuning
Table 3 compares our model’s ability to predict HSQC cross peaks at different training stages, highlighting the contributions of our training strategy. After pre-trained via MTT on the 1D NMR dataset, the model achieved the validation performance with mean absolute errors (MAEs) of 0.210 ppm for ${}^{1}\text{H}$ NMR prediction and 2.228 ppm for ${}^{13}\text{C}$ NMR prediction. This success can be attributed to MTT which allows the model to effectively learn atomic latent features as well as local structural information by simultaneously performing ${}^{1}\text{H}$ and ${}^{13}\text{C}$ NMR shift predictions. This helps us surpass the problem with limited annotated HSQC data. Interestingly, the pre-trained model was able to predict HSQC cross peaks with reasonable MAEs of 1.397 ppm and 2.822 ppm for ${}^{1}\text{H}$ and ${}^{13}\text{C}$ shifts, respectively. These relatively large MAEs are expected as the data distribution of the HSQC dataset (76.34% small molecules and 90.33% non-saccharides) differs significantly from that of the 1D NMR dataset (98.80% small molecules and 99.95% non-saccharides). In addition, the HSQC cross peaks involve interactions beyond simple pairings of 1D ${}^{13}\text{C}$ and ${}^{1}\text{H}$ shifts, requiring a deeper understanding of interactions between atoms. Finally, the frequent absence of solvent labels in the 1D NMR dataset prevents the model from learning solvent effects.
Table 3: MAEs for 2D NMR prediction at different stages of training
Training Stage 1H MAEs (ppm) 13C MAEs(ppm)
MTT 1.397 (0.085) 2.822 (0.092)
MTT+ IUL 0.163 (0.003) 2.035 (0.088)
Nevertheless, the pre-training via MTT offers a robust foundation for fine-tuning the model using IUL. With each IUL iteration, we observed a reduction in model errors. The performance improvement is more pronounced during the initial iterations and gradually diminishes. By the fifth iteration, the improvement became marginal, indicating the convergence of fine-tuning. Finally, the fine-tuned model achieves MAEs of 0.163 ppm and 2.035 ppm for ${}^{1}\text{H}$ and ${}^{13}\text{C}$ shifts, respectively. Throughout the IUL process, the model was trained to gain a more profound understanding of solvent effects and complex C–H interactions due to intricate molecular structures.
@@ 4 Conclusions and Discussions
In this study, we introduce a novel framework to develop machine learning techniques for predicting C–H cross peaks in HSQC spectra, The framework enables us to tackle two major challenges in this avenue. The first challenge is the scarcity of annotated HSQC data for training machine learning models. The second challenge is that collecting large volumes of annotated HSQC data is labor-intensive and requires highly trained personnel. In implementing our framework, we developed a model combining a GNN with a solvent encoder. The GNN is trained to generate atomic embeddings that encapsulate both the local and global chemical environments of each atom, which is crucial for accurate chemical shift predictions. The atomic embeddings are combined with the solvent embedding produced by the solvent encode, which allows our model to learn the influence of solvent on chemical shifts. The combined embeddings are mapped by MLP modules to HSQC chemical shifts. Our framework employs a two-stage transductive strategy to train the model while addressing the aforementioned challenges. In the first stage, we use a large amount of annotated 1D NMR data to pre-train the model via Multi-Task learning. This enables the model to adeptly grasp the intricate relationship between atomic interactions and NMR signals, laying a robust foundation for the subsequent stage. Next, the model is refined on a set of unlabelled HSQC spectra via Iterative Unsupervised Learning, enhancing the model’s capability in predicting and interpreting HSQC spectra. Our final model achieves MAEs of 0.163 ppm and 2.035 ppm for ${}^{1}\text{H}$ and ${}^{13}\text{C}$ shifts respectively, while accurately assigning cross peaks. It demonstrates a consistent performance across various molecular weight and saccharide categories, significantly outperforming the traditional methods, and shows convincing generalization capabilities to less represented samples from the training dataset. In the future, we plan to refine our model by developing 3D-GNN models that are able to consider 3D structural information such as spatial orientation and conformational flexibility. This enhancement should enable us to handle other 2D NMR spectra, such as Correlation Spectroscopy and Nuclear Overhauser Effect Spectroscopy, thus broadening its applicability and making a more substantial contribution to the field of chemical analysis.
@@ Acknowledgements
  • Scarselli et al. 2008 F. Scarselli, M. Gori, A. C. Tsoi, M. Hagenbuchner and G. Monfardini, IEEE transactions on neural networks, 2008, 20, 61–80.
  • Micheli 2009 A. Micheli, IEEE Transactions on Neural Networks, 2009, 20, 498–511.
  • Gilmer et al. 2017 J. Gilmer, S. S. Schoenholz, P. F. Riley, O. Vinyals and G. E. Dahl, International conference on machine learning, 2017, pp. 1263–1272.
  • Zhou et al. 2020 J. Zhou, G. Cui, S. Hu, Z. Zhang, C. Yang, Z. Liu, L. Wang, C. Li and M. Sun, AI open, 2020, 1, 57–81.
  • Reiser et al. 2022 P. Reiser, M. Neubert, A. Eberhard, L. Torresi, C. Zhou, C. Shao, H. Metni, C. van Hoesel, H. Schopmans, T. Sommer et al., Communications Materials, 2022, 3, 93.
  • Kuhn 1955 H. W. Kuhn, Naval research logistics quarterly, 1955, 2, 83–97.
  • Munkres 1957 J. Munkres, Journal of the society for industrial and applied mathematics, 1957, 5, 32–38.
  • Gold and Rangarajan 1996 S. Gold and A. Rangarajan, IEEE Transactions on pattern analysis and machine intelligence, 1996, 18, 377–388.
@@
  • 33 J-resonance, NMR Database, https://www.j-resonance.com/en/nmrdb/.
  • Foster et al. 2007 M. P. Foster, C. A. McElroy and C. D. Amero, Biochemistry, 2007, 46, 331–340.
@@
  • Brown et al. 2018 G. D. Brown, J. Bauer, H. M. Osborn and R. Kuemmerle, ACS omega, 2018, 3, 17957–17975.

### v3

@@ 3.4.1 Performance regarding molecular weight
Fig.  5: Model performance comparison between small, medium amd large molecules by molecular weight. The model performs equally well between small and medium molecules, with a marginally reduced precision for larger molecules.
@@
3.4.2 Performance on saccharides
Fig.  7: Examples of model performance on saccarides. The orange dots indicate the HSQC prediction by our model and the blue triangles indicates the observed signals from raw experiment data. For each example, the Mean Absolute Error (MAE) is shown in the lower right of the plot. Our model’s prediction is able to account for every signal that is observed from the experiments.
Saccharides, or carbohydrates, play critical roles in biological processes, serving as the primary source of energy for living organisms and functioning as essential structural components in cells and tissues.26, 27 Their ubiquitous presence and diverse functions underscore their importance in maintaining the delicate balance of life. From the glucose molecules that fuel cellular respiration to the intricate glycoproteins that mediate cellular recognition and signaling, carbohydrates form an indispensable cornerstone of biological systems.28, 29 Moreover, their dynamic interplay with proteins, lipids, and nucleic acids orchestrates a myriad of physiological processes, ranging from cell-cell communication to immune response modulation. 30 In essence, saccharides actively contribute to the complexity of life through their multifunctional roles and interactions within biological systems.
@@ 3.4.2 Performance on saccharides
However, elucidating the structures of saccharides presents a formidable challenge due to their inherent structural complexity and diversity. This complexity arises from the diverse arrangements of monosaccharide units, varied anomeric configurations, and variable glycosidic linkages. Additionally, carbohydrates often lack the requisite crystallinity, unlike the well-defined crystalline structures of small molecules or proteins, which hinders the acquisition of high-resolution X-ray diffraction data. Consequently, NMR spectroscopy, particularly through techniques such as HSQC, has emerged as an indispensable tool in unraveling the detailed structures of carbohydrates.31
@@
Our model demonstrates excellent performance in predicting HSQC spectra for saccharides, achieving a level of accuracy comparable to that observed for non-saccharide molecules, as illustrated in Figure 6. For non-saccharides, the HSQC prediction yields a MAE of 1.765 for ${}^{13}\text{C}$ shifts and 0.138 for ${}^{1}\text{H}$ shifts. For saccharides, the HSQC prediction yields a MAE of 1.808 for ${}^{13}\text{C}$ shifts and 0.163 for ${}^{1}\text{H}$ shifts. Furthermore, some examples are presented in Figure 7. These saccharide molecules feature multiple ring structures and numerous stereogenic centers, contributing to the intricate nature of their HSQC spectra. Despite these inherent challenges, our proposed model exhibits exceptional accuracy in predicting the spectra for these molecules. This robust performance underscores the model’s capacity to navigate the complexities associated with saccharide molecules, thereby emphasizing its versatility and effectiveness across various applications in the field.
3.5 Comparison to traditional tools
In organic chemistry, the simulation of HSQC spectra holds significant importance for analyzing experimental HSQC sepctra. It enables researchers to assign crosspeaks to C–H bonds in the target molecule. Traditional approaches, encompassing web-based tools like nmrdb.org 32, software solutions such as ChemDraw22, and Mestronova23, have long served as go-to resources for this task. In our investigation, we present two compounds (One is simple, the other is more complex.) for the comparison of classical methods with our innovative model.
Our model demonstrates clear advantages over traditional tools, as shown in Figure 8. For the simple molecule (a), traditional methods like ChemDraw, Mestronova, and nmrdb.org exhibit moderate performance, with MAEs for 13C shifts ranging from 1.35 to 3.47 ppm, and for 1H shifts ranging from 0.18 to 0.27 ppm. In contrast, our model’s predictions closely align with experimental observations, with MAEs of only 0.35 ppm for ${}^{13}\text{C}$ and 0.07 ppm for ${}^{1}\text{H}$. This disparity is even more pronounced for the larger molecule (b), where traditional methods show significantly larger MAEs ranging from 9.63 to 20.02 ppm for 13C shifts and from 0.74 to 1.19 ppm for 1H shifts. In contrast, our model achieves MAEs of only 0.67 ppm for ${}^{13}\text{C}$ shifts and 0.07 ppm for ${}^{1}\text{H}$ shifts.
Fig.  8: Performance comparison between proposed GNN model and traditional method. For each example (a) and (b), the observed signals from experiments are annotated by blue dots and the prediction by each method is highlighted in orange. The prediction error, measured by MAE, is shown in the bottom right on each plot. For a small molecule (a), all methods perform within reason but our model offers the most precise performance. For a large molecule (b), the advantage of our model becomes more evident, as the traditional methods become less accurate in analyzing complicated conformations.
@@ 4 Conclusions and Discussions
In this work, we introduced a novel approach for predicting HSQC spectra, aimed at addressing the challenges arising from significant computation cost, limited data availability and labor-intensive annotation. Our framework integrates a GNN and solvent impact to address these challenges effectively. The GNN facilitates the generation of atomic embeddings that encapsulate both local and global chemical environments surrounding each atom. These atomic embeddings capture intricate chemical properties and interactions that profoundly influence chemical shifts. Subsequently, multilayer perceptron (MLP) layers are employed to predict C–H shifts based on these refined atom embeddings. This approach effectively captures solvent effects, thereby enabling accurate predictions of chemical shifts.
For future directions, we aim to broaden the application of GNN-based NMR predictions to complex biological molecules, including proteins, nucleic acids, and large polysaccharides. Accurate prediction of NMR spectra for these molecules holds the potential to deepen our understanding of their structure-function relationships, dynamics, and interactions with other molecules. This advancement could significantly contribute to accelerating drug discovery and biochemical research efforts.
@@ Acknowledgements
  • 25 J-resonance, NMR Database, https://www.j-resonance.com/en/nmrdb/.
@@
  • Sun et al. 2016 L. Sun, D. R. Middleton, P. L. Wantuch, A. Ozdilek and F. Y. Avci, Glycobiology, 2016, 26, 1029–1040.